Maple Mount Fishery is a canning company in Astoria. The company uses a normal costing system in which factory overhead is applied on the basis of direct labor costs. Budgeted factory overhead for the year was $680,000, and management budgeted $320,000 of direct labor costs. During the year, the company incurred the following actual costs. Direct materials used Direct labor Factory overhead $382,000 313,000 650,700 Skipped The January 1 balances of inventory accounts are shown below Materials - all direct Work-in-process Finished goods $64,000 41,400 25, 600 The December 31 balances of these inventory accounts were ten percent lower than the balances at the beginning of the year. The adjusted cost of goods sold, after under or overapplied overhead, is: (Round your predetermined overhead rate to 1 decimal place.)Multiple Choice $1,352,300. $1,355,300
Answer #1

Overhead rate = 680/320 = 2.1 of direct labor cost

Total manufacturing cost = 382000+313000+(313000*2.1) = 1352300

Cost of goods manufactured = 1352300+41400-37260 = 1356440

Applied overhead = 313000*2.1 = 657300

Actual overhead = 650700

Overapplied overhead = 657300-650700 = 6600

Un adjusted cost of goods sold = 25600+1356440-23040 = 1359000

Adjusted cost of goods sold = 1359000-6600 = 1352400

So answer is e) $1352400